# E-chat: Emotion-sensitive Spoken Dialogue System with Large Language Models

### v1

@@ Abstract
This study focuses on emotion-sensitive spoken dialogue in human-machine speech interaction. With the advancement of Large Language Models (LLMs), dialogue systems can handle multimodal data, including audio. Recent models have enhanced the understanding of complex audio signals through the integration of various audio events. However, they are unable to generate appropriate responses based on emotional speech. To address this, we introduce the Emotional chat Model (E-chat), a novel spoken dialogue system capable of comprehending and responding to emotions conveyed from speech. This model leverages an emotion embedding extracted by a speech encoder, combined with LLMs, enabling it to respond according to different emotional contexts. Additionally, we introduce the E-chat200 dataset, designed explicitly for emotion-sensitive spoken dialogue. In various evaluation metrics, E-chat consistently outperforms baseline LLMs, demonstrating its potential in emotional comprehension and human-machine interaction.
Index Terms—  Large Language Model, emotional speech comprehension, spoken dialogue systems.
@@ 1 Introduction
Fig. 1: Emotion-senstive spoken dialogue scenario. <>absent<>< > denotes the emotion of the speech.
@@
In our evaluation process, we employ a comprehensive set of subjective and objective metrics to assess the response quality of E-chat. Our findings indicate that E-chat significantly outperforms baseline LLMs across objective metrics. In terms of subjective evaluations, E-chat also demonstrated superior performance, as reflected in its higher Mean Opinion Score (MOS), underscoring its exceptional capability to deliver emotionally nuanced responses.
@@ 2 Related work
Spoken Dialogue with Emotion. Recent studies have explored emotional responses within spoken dialogue systems. Some of these efforts focus on predicting emotion labels [20, 21], while others use these labels to guide the generation of response texts [22, 23]. ParalinGPT [24], by integrating text and speech modalities, more accurately simulates the linguistic content and paralinguistic attributes of spoken responses. In contrast, E-chat employs an end-to-end approach, enabling direct responses based on the emotional features in speech without the need for prior prediction of emotion or response labels.
@@ 3 Method
Fig. 2: Architecture schematic diagram of E-chat. <>absent<>< > denotes the emotion of the speech. ()()( ) denotes the translation of Chinese sentences. On the left is the architecture of E-chat. On the right are the corresponding data samples and prompts for the two-stage training.
@@ 3.1 E-chat Architecture
The speech encoder extracts both speech and emotion features, with the former transformed into the LLM decoder’s textual feature space via a connection module. This transformation is crucial as it ensures that the speech features are compatible with the LLM’s processing capabilities. The concatenation of emotion features with speech features is an important step to enrich the decoder input, providing a more holistic view of the speaker’s emotional state. Ultimately, the LLM decoder, armed with these comprehensive features, is better equipped to generate responses that are not only contextually relevant but also emotionally attuned, thereby elevating the conversational quality of the model.
@@
LLM Decoder. This study uses the Chinese version of LLaMA-7B [5], Atom-7B  [^2] https://huggingface.co/FlagAlpha/Atom-7B, as the LLM decoder for training. Based on the LLaMA structure, Atom-7B undergoes further pre-training with extensive Chinese data, thereby excelling in various Chinese reasoning and generation tasks. Atom is a model consisting of a 32-layer Transformer decoder with a total of 7.7B parameters. To prevent catastrophic forgetting in the LLM, the parameters of the LLM are frozen during the entire training phase.
@@ 3.2 Two-Stage Training
4 Echat-200h
For optimal training, E-chat requires a dataset designed for emotion-sensitive spoken dialogue. Recognizing the inadequacy of existing datasets for E-chat’s distinct requirements, we develop E-chat200 dataset. Each entry comprises a tuple of (question text, response, emotion, speech). Within this structure, the speech serve as inputs to the model, while the emotion and response act as labels.
@@ 4 Echat-200h
Table 1 presents the statistical data of the E-chat200 dataset. E-chat200 dataset encompasses a total of 184,000 tuples of (question text, response, emotion, speech), with emotions categorized into five types: cheerful, fearful, angry, sad, and neutral, covering a range of ten different speakers. This dataset has been pivotal for the successful training of E-chat.
Upon constructing E-chat200 dataset, we randomly select 10k entries for testing, 10k entries for validation, and the remaining 164k entries for training.
@@
Cheerful 62 74 37
Fearful 25 20 10
Angry 33 38 19
Sad 30 34 17
Neutral 34 34 17
Sum 184 200 100
@@ 5.2 Evaluation Metric
Fig. 3: Examples of E-chat. Each sentence is manually recorded using an Android phone. <>absent<>< > denotes the emotion of the speech. ()()( ) denotes the translation of Chinese sentences. On the left are three samples, on the right are the corresponding questions and responses for other emotions.
We design objective and subjective evaluation methods based on the task of speech emotion caption [28].
@@
We compute SIM between E-chat and GPT-3.5-Turbo, which inputs entirely correct emotion labels and question text to establish a reference topline. To demonstrate E-chat’s emotional understanding, we compute SIM between Atom and GPT-3.5-Turbo and then compare it with E-chat, where Atom processes question text but without emotion labels. This comparison, using GPT-3.5’s responses as a topline, aims to highlight E-chat’s advanced capability in interpreting and responding to emotion in speech.
@@ 5.3 Results
Objective Results. E-chat model demonstrates superior emotion and speech understanding capabilities when compared with Atom. As shown in Table  2, in SIM, E-chat’s responses aligned closely with the responses provided by GPT-3.5, indicating its effectiveness in understanding and expressing emotions. Compared to the Atom model, E-chat exhibit a clear advantage in handling emotional queries (i.e., 72.70 vs 64.77 SIM score). Similar results are found for E-chat on BLEU scores.
Subjective Results. In the MOS scoring, E-chat receives high marks for the naturalness and accuracy of its emotional expressions. As shown in Table 2, the quality of responses generated by E-chat significantly surpasses that of Atom (i.e., 3.53 vs 2.71 MOS score) and is on par with emotion GPT-3.5 annotations (i.e., 3.53 vs 3.71 MOS score). This reflects E-chat’s advancement and practicality in emotion comprehensive.
@@
Metric E-chat stage-2 Atom-7B GPT-3.5 *
$\text{BLEU}_{1}$ 26.12 21.76 -
$\text{BLEU}_{4}$ 0.52 0.44 -
SIM 72.70 64.77 -
MOS 3.53 2.71 3.71
Examples. In Fig. 3, we present some examples of E-chat’s application in spoken dialogue system. To test E-chat’s performance on recorded voice, each sentence is manually recorded using an Android phone. The results demonstrate E-chat’s accurate understanding of all questions, attributed to its extensive use of ASR data in stage-1, ensuring robustness to recorded voice. Additionally, each question was recorded in two different emotional tones to verify E-chat’s ability to respond to various emotions. The outcomes confirm E-chat’s effective emotion recognition and appropriate response, aligning with our expectations.
@@ 5.4 Analysis and Discussion
Two-stage Training. In fact, during our research, we initially attempt to train the entire E-chat model solely using stage-2. The results indicate that the model only learned emotional and subject information within sentences. Consequently, it is crucial to determine whether the model developed in stage-1 could effectively transform the feature space. To verify this, we analyze the performance of the stage 1 model. We employ results from ASR to test the model obtained from stage-1, as shown in Table 3. The findings demonstrate that our model exhibits competitive and, in some cases, superior performance compared to Whisper-large-v2 [12] on both test sets of WenetSpeech and the test sets of AISHELL-1 and AISHELL-2. This suggests that the stage-1 model is capable of adeptly transforming speech embeddings into a feature space suitable for LLM input, laying a crucial foundation for the training in stage-2.
@@
E-chat stage-1 3.87 6.07 12.93 17.86
Speech Emotion Recognition. A key objective of E-chat is to provide appropriate responses based on emotion, making emotion recognition accuracy a vital metric. We assess this aspect using E-chat200 test set. The results reveal that E-chat achieves an accuracy rate of 74.1% in emotion recognition, indicating a commendable capability. Compared to other speech emotion recognition systems [29], E-chat is similar in relevant metrics, demonstrating reliable emotional comprehension. Furthermore, the use of soft emotion embeddings for E-chat input to the decoder enhances the accuracy of appropriate responses in practical applications. These findings not only validate E-chat’s effectiveness in comprehending various emotions but also lay a solid foundation for future emotionally intelligent human-machine interactions.
@@ 6 Conclusion
This study addresses the limitation of existing LLMs in generating appropriate responses to emotional speech. We introduce E-chat model, enhanced by E-chat200 dataset, specifically designed for emotion-sensitive spoken dialogue. Our evaluations demonstrate E-chat’s capability to discern and respond to various emotional cues in speech, marking a significant advancement in this area. However, the model currently faces limitations in handling audio with mixed emotions. Future research will focus on improving this aspect, thereby refining E-chat’s applicability in more complex emotional speech scenarios.

### v3

@@ Abstract
This study focuses on emotion-sensitive spoken dialogue in human-machine speech interaction. With the advancement of Large Language Models (LLMs), dialogue systems can handle multimodal data, including audio. Recent models have enhanced the understanding of complex audio signals through the integration of various audio events. However, they are unable to generate appropriate responses based on emotional speech. To address this, we introduce the Emotional chat Model (E-chat), a novel spoken dialogue system capable of comprehending and responding to emotions conveyed from speech. This model leverages an emotion embedding extracted by a speech encoder, combined with LLMs, enabling it to respond according to different emotional contexts. Additionally, we introduce the E-chat200 dataset, designed explicitly for emotion-sensitive spoken dialogue. In various evaluation metrics, E-chat consistently outperforms baseline model, demonstrating its potential in emotional comprehension and human-machine interaction  [^1] E-chat samples: https://anonymous-echat.github.io/E-chat/.
Index Terms: large language model, emotional speech comprehension, spoken dialogue systems
@@ 1 Introduction
Figure 1: Emotion-senstive spoken dialogue scenario. <>absent<>< > denotes the emotion of the speech.
@@
In our evaluation process, we employ a comprehensive set of subjective and objective metrics to assess the response quality of E-chat. Our findings indicate that E-chat outperforms baseline model across objective metrics. In terms of subjective evaluations, E-chat also demonstrated superior performance, as reflected in its higher Mean Opinion Score (MOS), underscoring its exceptional capability to deliver emotionally nuanced responses.
@@ 2 Related work
Spoken Dialogue with Emotion. Recent studies have explored emotional responses within spoken dialogue systems. Some of these efforts focus on predicting emotion labels [20, 21], while others use these labels to guide the generation of response texts [22, 23]. ParalinGPT [24] integrates text and speech modalities to more accurately simulate both the linguistic content and paralinguistic attributes of spoken responses. However, ParalinGPT uses an ASR model to obtain text and a speech encoder to extract emotion embeddings. In contrast, E-chat employs an end-to-end approach, generating direct responses based on the textual and emotional features in speech without needing prior prediction of emotion or response labels.
@@ 3 Method
Figure 2: Architecture schematic diagram of E-chat. <>absent<>< > denotes the emotion of the speech. ()()( ) denotes the translation of Chinese sentences. On the left is the architecture of E-chat. On the right are the corresponding data samples and prompts for the two-stage training.
@@ 3.1 E-chat Architecture
The E-chat model, depicted in Fig.2, employs a speech encoder to extract speech and emotion features. These are then integrated, transforming speech features for compatibility with the LLM decoder, and enriching inputs with emotional features. This process enables the LLM decoder to produce responses that are both relevant and emotionally resonant, improving conversation quality.
@@
LLM Decoder. This study uses the Chinese LLM, Baichuan2-7B-Chat  [^3] https://huggingface.co/baichuan-inc/Baichuan2-7B-Chat, as the LLM decoder for training. Based on the LLaMA structure, Baichuan2-7B-Chat undergoes pre-training with extensive Chinese data, thereby excelling in various Chinese reasoning and generation tasks. Baichuan2-7B-Chat is a model consisting of a 32-layer Transformer decoder with a total of 7.7B parameters. To prevent catastrophic forgetting in the LLM, the parameters of the LLM are frozen during the entire training phase.
@@ 3.2 Two-Stage Training
4 Echat-200
For optimal training, E-chat requires a dataset designed for emotion-sensitive spoken dialogue. Recognizing the inadequacy of existing datasets for E-chat’s distinct requirements, we develop E-chat200 dataset. Each entry comprises a tuple of (question text, response, emotion, question speech). Within this structure, the question speech serve as inputs to the model, while the emotion and response act as labels.
@@ 4 Echat-200
Table 1 presents the statistical data of the E-chat200 dataset. E-chat200 dataset encompasses a total of 178k tuples of (question text, response, emotion, speech), with emotions categorized into five types: cheerful, fearful, angry, sad, and neutral, covering a range of ten different speakers. This dataset has been pivotal for the successful training of E-chat.
Upon constructing E-chat200 dataset, we randomly select 23k entries for testing, 12k entries for validation, and the remaining 143k entries for training. We will open source this dataset after the paper is accepted.
@@
Cheerful 62 73 38
Fearful 25 21 11
Angry 30 33 17
Sad 33 38 20
Neutral 28 28 14
Sum 178 193 100
@@ 5.2 Evaluation Metric
Figure 3: Examples of E-chat. Each sentence is manually recorded using an Android phone. <>absent<>< > denotes the emotion of the speech. ()()( ) denotes the translation of Chinese sentences. On the left are three samples, on the right are the corresponding questions and responses for other emotions.
We designed both objective and subjective evaluation methods based on the task of speech emotion captioning [28]. We implement ParalinGPT [24] as a baseline system, where the ASR model used is Whisper-large-v2, and the LLM employed is Baichuan2-7B-Chat.
@@
We compute SIM between E-chat and GPT-3.5-Turbo, which inputs entirely correct emotion labels and question text to establish a reference topline. This comparison, using GPT-3.5’s responses as a topline, aims to highlight E-chat’s advanced capability in interpreting and responding to emotion in speech.
@@ 5.3 Results
Objective Results. E-chat model demonstrates superior emotion and speech understanding capabilities when compared with ParalinGPT. As shown in Table 2, in SIM, E-chat’s responses aligned closely with the responses provided by GPT-3.5, indicating its effectiveness in understanding and expressing emotions. Compared to the ParalinGPT, E-chat exhibit a clear advantage in handling emotional queries (i.e., 74.83 vs 65.84 SIM score). Similar results are found for E-chat on BLEU scores. This may be due to the fact that ParalinGPT is using the results of the ASR model as input to the LLM, which leads to error accumulation.
Subjective Results. In the MOS scoring, E-chat receives high marks for the naturalness and accuracy of its emotional expressions. As shown in Table 2, the quality of responses generated by E-chat surpasses that of ParalinGPT (3.56 vs. 3.37 MOS score) and is comparable to emotion GPT-3.5 annotations (3.56 vs. 3.71 MOS score). This demonstrates E-chat’s advancement and practicality in comprehensive emotion recognition.
@@
Metric E-chat ParalinGPT [24] GPT-3.5 *
$\text{BLEU}_{1}$ 31.07 20.19 100
$\text{BLEU}_{4}$ 9.1 4.42 100
SIM 74.83 65.84 100
MOS 3.56 3.37 3.71
Examples. In Fig. 3, we present some examples of E-chat’s application in spoken dialogue system. To test E-chat’s performance on recorded voice, each sentence is manually recorded using an Android phone. The results demonstrate E-chat’s accurate understanding of all questions, attributed to its extensive use of ASR data in stage-1, ensuring robustness to recorded voice. Additionally, each question is recorded in two different emotional tones to verify E-chat’s ability to respond to various emotions. The outcomes confirm E-chat’s effective emotion recognition and appropriate response, aligning with our expectations.
@@ 5.4 Analysis and Discussion
Two-stage Training. In fact, during our research, we initially attempt to train the entire E-chat model solely using stage-2. The results indicate that the model only learned emotional and subject information within sentences. Consequently, it is crucial to determine whether the model developed in stage-1 could effectively transform the feature space. To verify this, we analyze the performance of the stage-1 model. We employ results from ASR to test the model obtained from stage-1, as shown in Table 3. The findings demonstrate that our model exhibits superior performance compared to Whisper-large-v2 [12] on both test sets of WenetSpeech and the test sets of AISHELL-1 and AISHELL-2. This suggests that the stage-1 model is capable of adeptly transforming speech embeddings into a feature space suitable for LLM input, laying a crucial foundation for the training in stage-2.
@@
E-chat stage-1 1.31 3.48 6.55 6.84
Speech Emotion Recognition. A key objective of E-chat is to provide appropriate responses based on emotion, making emotion recognition accuracy a vital metric. We assess this aspect using E-chat200 test set. The results reveal that E-chat achieves an accuracy rate of 73.6% in emotion recognition, indicating a commendable capability. Compared to other speech emotion recognition systems [29], E-chat is similar in relevant metrics, demonstrating reliable emotional comprehension. Furthermore, the use of soft emotion embeddings for E-chat input to the decoder enhances the accuracy of appropriate responses in practical applications. These findings not only validate E-chat’s effectiveness in comprehending various emotions but also lay a solid foundation for future emotionally intelligent human-machine interactions.
@@ 6 Conclusion
This study addresses the limitations of existing large language models in generating appropriate responses to emotional speech. We introduce the E-chat model, enhanced by the E-chat200 dataset, which is specifically designed for emotion-sensitive spoken dialogue. Our evaluations demonstrate E-chat’s capability to discern and respond to various emotional cues in speech, marking a significant advancement in this field. However, the model currently only produces text, a limitation that can be addressed by integrating a text-to-speech model. In the future, we plan to explore end-to-end emotion-sensitive speech-to-speech models.